# The modelling of the action potentials in myelinated nerve fibres

### v2

@@ 3.2 Saltatory conduction mechanism hypothesis for myelinated axon
It is known that the myelination of the axon increases the propagation velocity of the AP and the prevalent explanation in earlier studies of this phenomenon is the saltatory conduction mechanism [26, 15, 55, 19, 32]. As the capacity of the axon changes significantly between the myelinated and unmyelinated sections, this causes, in a nutshell, the electrical signal to “jump" between the nodes of Ranvier propagating faster than it would in the case of an unmyelinated axon. The saltatory conduction hypothesis is briefly summarised as follows.
@@
In practice, the saltatory conduction hypothesis, although difficult to verify experimentally as measuring sharply localised AP (nodes of Raniver are typically around 1 µ long) is measured [19]. What is measured in classical experiments (for example, [6, 56, 51]), is AP over some section of the axon (i.e, including the signal from both myelinated and non-myelinated sections, moreover most classical experiments are done on non-myelinated axons). However, the proposed hypothesis appears to be a plausible (and widely used in earlier studies) starting point to explain why the AP is faster in myelinated axons and we opt to use it as well. It should be noted that in effect we use interpretation which is partly inspired from the continuum mechanics [57] where we interpret the signal coming from the model not as the evolution of AP in time at a fixed point on axon as is typical but rather a combined signal one might get from a “unit cell" which contains a node of Ranvier and a myelinated section next to the node, based on assumption that characteristic wavelengths of the signal are much larger than the underlying scale of the micro-structure (nodes of Ranvier in µ range and myelinated sections typically in hundreds of µ in length [29, 13]). In comparison, the first harmonics (when looking at the Fourier spectrum of the signal) of the AP in space are from to tens of (depending on the duration and velocity of the signal) [58, 59].
@@
Here , , and is the length of the node of Ranvier. The parameter (16) governs the saltatory conduction velocity between adjacent nodes of Ranvier under the simplifying assumptions done above and is a reasonable starting point. It is worth noting that the philosophy behind constructing eq. (15) is one of the typical approaches how the HH equations (which in its ODE form describes signal evolution in time at a fixed spatial point on axon) are used to describe travelling AP signal – meaning that one constructs a pseudo-numerical scheme where adjacent nodes interact and can propagate the signal that way along the axon.
@@ 4 Model for AP propagation on myelinated axon based on Lieberstein model
Taking the elementary form of Maxwell equations combined with the ideas proposed by Lieberstein in eqs. (8) and (9) as a starting point, we proceed to modify these governing equations to include the effect of myelination on the AP signal propagation. The simplified geometry of a myelinated axon is shown in Fig. 5.
@@ 4.1 Phenomena and descriptions
When we modify the Lieberstein model [23] to account for the effect of myelination on a nerve fibre, we consider the following phenomena:
@@
| Table 3: AP velocity against -ratio. | |||||||
| : | 0 | 25 | 50 | 75 | 100 | 125 | 150 |
| µ | 1 | 5.3 | 7.5 | 9 | 10.5 | 11.7 | 12.8 |
| µ | 1.9 | 9.9 | 13.9 | 17 | 19.6 | 21.8 | 23.9 |
| µ | 2.5 | 13 | 18.2 | 22.3 | 25.7 | 28.6 | 31.4 |
| µ | 3.2 | 16.3 | 22.8 | 27.8 | 32 | 35.8 | 39.2 |
| µ | 3.8 | 19 | 26.7 | 32.7 | 37.6 | 42 | 46 |
| : | 175 | 200 | 225 | 250 | 275 | 300 | 325 |
| µ | 13.8 | 14.8 | 15.7 | 16.5 | 17.3 | 18 | 18.8 |
| µ | 25.9 | 27.6 | 29.3 | 30.9 | 32.4 | 33.8 | 35.2 |
| µ | 33.9 | 36.2 | 38.4 | 40.5 | 42.4 | 44.3 | 46.1 |
| µ | 42.3 | 45.2 | 47.9 | 50.5 | 53 | 55.3 | 57.7 |
| µ | 49.7 | 53 | 56.3 | 59.3 | 62.2 | 65 | 67.7 |
In eq. (17) parameter (describing -ratio) describes the average length of the myelinated section divided by the average length of the node of Ranvier (see Fig. 5). It affects the quantity which is the current along the axis of the axon. Parameter is a phenomenological coefficient which determines conduction velocity between adjacent nodes of Ranvier (generalised from eq. (16)). Here parameter includes myelin geometry perpendicular to the axon (related to -ratio). As noted earlier, parameter is not similar to parameter (see eq. (16)) and is a generalised quantity. We assume parameter to be between 0 and 1: if it is equal to 0, then the current can not propagate between the adjacent nodes (nodes of Ranvier are isolated from each other) and if it is equal to 1, then the myelinated sections are almost perfect conductors (i.e., the system obeys Ohm’s Law) and adjacent nodes react to the changes in a given node almost immediately. Parameter could be considered like a generalised dimensionless relative velocity potentially containing all the physical effects (not only resistance but also diffusive, capacitative, inductance effects, the influence of cytoskeleton or even damage or pathological effects, etc) as it is chosen here. One should also stress that in the following example, we take for the sake of simplicity, however, it would be logical that as the length of myelinated sections increases then at some point the parameter would start to decrease. Using the same parameters as in the classical paper by Hodgkin and Huxley [1], a numerical example using parameter value (assuming 1 µ nodes of Ranvier, 50 µ length for myelinated sections and taking for the numerical example) follows (see Fig. 6). Choosing (which is an upper limit case) means that we are making here an assumption that saltatory conduction velocity is much faster than signal propagation velocity in an unmyelinated axon. We have to emphasize here that the “unit cell" we are modelling in the computational node contains both the node of Ranvier and the myelinated section adjacent to it, i.e., the parameters are a mix of “pure” Lieberstein model for unmyelinated axon and myelination effects like in [26].
@@
The AP propagation velocity as a function of the -ratio at a fixed axon radius ( µ, µ, µ, µ, and µ) is depicted in Fig. 7. The case corresponds to an unmyelinated axon. It should be noted that here we are assuming that the thickness of the myelin (related to g-ratio, which is included through coefficient (16)) is the same even if -ratio is varied. The upper limit case () for the conduction velocity is used. It should be noted that the dependence of the AP velocity on the axon radius is still there even in the myelinated case. The larger the radius of the axon, the faster the velocity of the AP increases as the -ratio is increasing.
4.2 On AP energy in myelinated axon
When analysing the propagation of the AP, several forms of energy could be considered: capacitative, heat, inductive and even chemical. What follows is a short discussion on the AP energy based on our studies in addition to the previous analysis [21, 38, 37, 60, 61].
As a first rough estimate of just one part for the energy of AP, we can take it as the energy contained in an electric field on a capacitor. Taking the simplified axon geometry as a cylinder with a charge difference across the membrane one can make a naive simplification that the capacitance of such a cylinder segment is the same as a classical plate capacitor
where is the electric constant or vacuum permittivity, is the dielectric constant (roughly 7 for lipid bi-layer), is area (for cylinder wall where is the radius and is the length) and is the thickness of insulating layer between the plates. As an example, for a rough estimate let us take µ, and . Inserting the noted values into the (22), we get capacitance of roughly (per of axon). However, it must be noted, that the value of the dielectric constant can vary across protein internal structure and its changes [62].
The energy stored in the electric field in a charged parallel plate capacitor can be expressed as
where is capacitance and is potential difference. Assuming , we get a rough estimate of the energy as per cm.
This is a rough estimate for an unmyelinated axon (i.e., the node of Ranvier) because any geometry variations or possible changes of the dielectric constant during AP generation are neglected. Also, it must be noted that the electric field is quite significant in the unmyelinated sections (approximately potential difference across only roughly distance yields an electric field of ).
Now if we add into our consideration the myelination it seems reasonable to assume that energy must be conserved, i.e., when the AP signal is “jumping" between the nodes of Ranvier the energy of the pulse is roughly the same per cm as we estimated for the node of Ranvier where the main difference is that the thickness of the insulating layer is much larger. For example, let us take µ. In such a case the capacitance estimate drops to roughly per cm if all other parameters remain the same and the average electric field in the insulating layer drops to about for a myelination thickness.
However, one should note that in the myelinated case, the signal is substantially faster and as a result the AP is spread over a significantly longer section of the axon in the myelinated case. One can argue that if considering the “total" AP signal energy (over its whole duration, at a fixed spatial point or at a fixed time over its whole spatial distribution) it could be generalised (drawing inspiration from Poynting theorem in 1D case in classical electrodynamics [63]) to consider the integral instead of just in (23). Considering this idea we can combine eqs. (22) and (23) replacing the with its integral over half-space can write
as a rough estimate of the total energy of the electric field for AP in space or in time, respectively. Note that we have assumed here that only a single AP pulse exists in the half-space or within the given timeframe for eq. (24). It should be emphasized, that unlike classical capacitor energy expression (23) that yields estimate in joules the proposed generalised expressions (24) yield estimates on joule meters and seconds, respectively.
On the other hand, considering only the question of capacitor energy is only one aspect of the AP energy. One could ask how the inclusion of inductivity affects the calculation of AP energy? As a first approximation, one could be tempted to consider the energy in transmission line (1), (2) including inductivity (see Fig. 8 for a scheme of a unit cell) as opposed to just capacitor energy as it is usually done in studies based on the HH equation. However, authors of the present paper remain somewhat sceptical towards such an approach as it is not entirely clear if considering magnetic field energy on such a small scale in the electrically conductive environment inside an axon would yield a reasonable estimate, even if we insist on including the noted inductivity. After all, the model proposed here does not include the influence of the cytoskeleton explicitly (although its influence on the AP propagation is taken into account indirectly through parameter ).
Another possibility is to approach the question of the energy of AP from the perspective of power. It is known that power is , where is the current in a circuit and is the potential difference. Considering the power we could separate it into two generalised components, first, across the membrane at a fixed spatial point where is the current across the membrane (taken similar to the HH equations here) and is the AP amplitude at the given spatial point at given time (about 100 at the peak) and second, along the axon at the fixed time where is current along the axis of the axon and is the potential gradient along the axis of the axon. The total power of the electrical pulse would be then .
@@ 5 Discussion and summary
We started with the elementary form of Maxwell equations (1), (2) and drawing inspiration from the classical HH paper [1] used the total membrane current density (6) to derive the model equations for unmyelinated axon similar to the Lieberstein model [23]. While Lieberstein opted to go a step further by moving into a moving frame of reference we opted to stay with the form which is closer to the Maxwell equations for a transmission line eqs. (8), (9) and used the work of Lieberstein as a source of inspiration for handling the question of inductance in the context of signal propagation along the nerve axon. Using the parameters from the literature we investigated briefly the solutions of the noted model for the unmyelinated axon and demonstrated that the behaviour of the solutions is in the physiologically plausible range and the key characteristics of the nervous signalling are fulfilled. These are: (i) the annihilation of AP signals during a head-on collision, (ii) the existence of activation threshold, and (iii) the refraction period after signal passing. In the parameter range considered, we have calculated the AP signal propagation velocity from 0.5 up to about 4.4 for the unmyelinated axon. The key difference between the classical HH equations and the Lieberstein-inspired model used here is that the mechanism for signal propagation along the axon emerges like a wave as a consequence of opting to keep the inductivity . While, indeed, there exist variations of the classical HH equations which support AP signal propagation where the equations are written in the form of PDE instead of the usual ODE form (which describes signal evolution in time at a fixed spatial point). In these, normally, the potential-gradient-type member responsible for propagating the signal along the axon is not as clearly defined from the viewpoint of physics (usually some kind of abstracted diffusion-type process is used). This is essential later in the paper for the purpose of clearer physical interpretation as we make use of the model based on the elementary form of Maxwell equations which is further modified to include the influence of myelination on the signal propagating along the axon.
@@
We briefly explained the saltatory conduction mechanism hypothesis as we prepared to modify the governing equations to take into account the effect of myelination on the propagation of the AP along the axon. The model (8), (9) based on Maxwell equations for a transmission line is then modified to include the influence of myelination. In addition to the g-ratio normally considered in the earlier studies (taken into account indirectly through coefficient ) which takes into account the myelination geometry perpendicular to the axon we introduce the so-called “myelination-ratio" or -ratio which describes the influence of myelin distribution on the signal propagation in the direction of the axis of the axon. The numerical example (see Fig. 7), using parameters from the earlier studies, demonstrates physiologically plausible behaviour for the model. The equations are reduced to the description of the unmyelinated axon if the length of the myelinated sections along the axon is taken as zero. Under the considered parameter combinations we can observe the AP propagation velocities up to 67.7 (the signal propagation velocity range for myelinated axons is given as roughly 10 to 100 in the earlier studies [29, 30]).
@@
It is important to emphasize that the proposed continuum-based model is philosophically similar to how the transmission line equations are composed. The ‘unit-cell’ in the context of the myelinated axon in the model is composed of the node of Ranvier and the myelinated section next to it. This is opposed to the alternative approach where the processes in nodes of Ranvier are described by the classical HH equations while myelinated sections are handled separately either through some numerical scheme or by an alternative model coupled with the HH equations in the node of Ranvier through some mechanism. Having a relatively simple pair of PDEs which are connected to the Maxwell equations for anything involving the movement of charges in an environment could be considered superior to investigating causal connections and making predictions than something that is not as clearly connected to the basic physical considerations.
Finally, we briefly discuss the question of the energy of the AP pulse in both myelinated and unmyelinated axons, making an argument that while the normally used capacitor-energy-based energy estimate is a reasonable starting point, something more could be needed. Speculating that perhaps instead of considering capacitor energy as the energy of an AP signal it could be better to consider the power of the signal (either in space, at a fixed time or in time, at a fixed spatial point).
@@ Acknowledgements
This research was supported by the Estonian Research Council (PRG 1227). Jüri Engelbrecht acknowledges the support from the Estonian Academy of Sciences. The authors thank Prof. Antoine Jerusalem from the University of Oxford for helpful comments during the preparation of the present paper. The authors thank the anonymous reviewers for numerous suggestions for improvements and helpful comments.
Author contributions Kert Tamm contributed by writing manuscript text and composing figures, performing numerical simulations and analysis of the results. Tanel Peets contributed analysis for comparing the modified Lieberstein equation against the classical Hodgkin-Huxley equations. All other work (like novel ideas presented, physical interpretation of the results, proofreading, etc) is a collaborative effort equally contributed by all authors.

### v3

@@ 9.2 Saltatory conduction mechanism hypothesis for myelinated axon
It is known that the myelination of the axon increases the propagation velocity of the AP and the prevalent explanation in earlier studies of this phenomenon is the saltatory conduction mechanism [3, 15, 50, 25, 18]. As the capacity of the axon changes significantly between the myelinated and unmyelinated sections, this causes, in a nutshell, the electrical signal to “jump” between the nodes of Ranvier propagating faster than it would in the case of an unmyelinated axon. The saltatory conduction hypothesis is briefly summarised as follows.
@@
In practice, the saltatory conduction hypothesis, although difficult to verify experimentally as measuring sharply localised AP (nodes of Raniver are typically around 1 µ long) is measured [25]. What is measured in classical experiments (for example, [26, 53, 24]), is AP over some section of the axon (i.e, including the signal from both myelinated and non-myelinated sections, moreover most classical experiments are done on non-myelinated axons). However, the proposed hypothesis appears to be a plausible (and widely used in earlier studies) starting point to explain why the AP is faster in myelinated axons and we opt to use it as well. It should be noted that in effect we use interpretation which is partly inspired from the continuum mechanics [47] where we interpret the signal coming from the model not as the evolution of AP in time at a fixed point on axon as is typical but rather a combined signal one might get from a “unit cell” which contains a node of Ranvier and a myelinated section next to the node, based on assumption that characteristic wavelengths of the signal are much larger than the underlying scale of the micro-structure (nodes of Ranvier in µ range and myelinated sections typically in hundreds of µ in length [33, 7]). In comparison, the first harmonics (when looking at the Fourier spectrum of the signal) of the AP in space are from to tens of (depending on the duration and velocity of the signal) [48, 12].
@@
Here , , and is the length of the node of Ranvier. Parameter (18) governs the saltatory conduction velocity between adjacent nodes of Ranvier under the simplifying assumptions done above and is a reasonable starting point. It is worth noting that the philosophy behind constructing eq. (17) is one of the typical approaches how the HH equations (which in its ODE form describes signal evolution in time at a fixed spatial point on axon) are used to describe travelling AP signal – meaning that one constructs a pseudo-numerical scheme where adjacent nodes interact and the signal can propagate that way along the axon.
@@ 10 Model for AP propagation on myelinated axon based on Lieberstein model
Taking the elementary form of Maxwell equations combined with the ideas proposed by Lieberstein [31] in eqs. (13) and (14) as a starting point, we proceed to modify these governing equations to include the effect of myelination on the AP signal propagation. The simplified geometry of a myelinated axon is shown in Fig. 9.
10.1 Parameters for numerical example
The following parameters are used: (number of spatial nodes), (time in [ms]), µ (membrane capacitance), µ (axon radius), (axoplasm resistance) while the parameters related to ion channel dynamics are the same as taken in [23]: and while (initial values for parameters in HH equations at ). We take in the following example the (resistance of an axon per unit length). We take (inductivity) for the numerical example (based on observed AP velocity for the HH model [23]). Initially we generate a narrow bell-shaped pulse (“spark”) for the in the middle of the space domain which generates the propagating AP. In spatial units, the length of the computation node () is 92 while the width of the computational domain in the space (from to ) is 24. For the sake of readability, the additional equations and values of physical quantities used in numerical simulation
are collected in Table 1 following [23] and [31]:
| Table 1: Parameters for finding example solutions for (13) and (14). | ||
It should be stressed, that the chosen parameters do not represent any particular model nerve as some of the parameters are varied (for example, axon radius) in a range that is outside what is typical for the experimentally observed (for the giant axon of the squid) or even taken as a rough estimate (like specific inductivity L). Most of the parameters are taken from [23] corresponding to giant axon of the squid at about while inductivity is chosen so that the AP signal would have propagation velocity of 21.2 if the axon radius is 238 [µ.
We use pseudospectral method (PSM) (see Appendix A in [11]) and periodic boundary conditions demonstrating the evolution of solutions for model equations (13) and (14). For initial condition we generate a narrow localised pulse for AP in the middle of the 1D space domain at time :
where [mV] is amplitude of the pulse, is the width of the pulse, is number of 2 sections in space in space and shifts the hyperbolic secant square function to the middle of the space domain (as it is defined around 0 but we have opted to integrate from 0 to 2 for the following numerical example) and take the other quantities initially as zero (at rest). Briefly, the main point of the pseudospectral method is that the discrete Fourier transform (DFT) based (PSM) (see also [17]) can be used to represent variable in the Fourier space as
reducing, therefore, the partial differential equation (PDE) to an ordinary differential equation (ODE) and then to use standard ODE solvers for integration in time. For integration in time, the model equations are rewritten as a system of first-order ODE’s and a standard numerical integrator is applied. In the numerical examples given in the present paper the ODEPACK FORTRAN code (see [22]) ODE solver is used through its NumPy implementation. Handling of the data and initialization of the variables is done in Python by making use of the package SciPy (see [27]) and the numerical results are analysed and visualised in the Matlab environment. Example solution can be seen in Fig. 10.
@@ 10.2 Phenomena and descriptions
We modify the Lieberstein model [31] to account for the effect of myelination on a nerve fibre, and consider the following phenomena:
@@
In eq. (22) parameter (describing -ratio) describes the average length of the myelinated section divided by the average length of the node of Ranvier (see Fig. 9). It affects the quantity which is the current along the axis of the axon. Parameter is a phenomenological coefficient which determines conduction velocity between adjacent nodes of Ranvier (generalised from eq. (18)). Here parameter includes myelin geometry perpendicular to the axon (related to -ratio). As noted earlier, parameter is not exactly the same as parameter (see eq. (18)) and is a generalised quantity. We assume parameter to be between 0 and 1: if it is equal to 0, then the current can not propagate between the adjacent nodes (nodes of Ranvier are isolated from each other) and if it is equal to 1, then the myelinated sections are almost perfect conductors (i.e., the system obeys Ohm’s Law) and adjacent nodes react to the changes in a given node almost immediately. Term could be considered like a generalised dimensionless relative velocity potentially containing all the physical effects (not only resistance but also diffusive, capacitative, inductance effects, the influence of cytoskeleton or even damage or pathological effects, etc) as it is chosen here. One should also stress that in the following example, we take for the sake of simplicity, however, it would be logical that as the length of myelinated sections increases then at some point the parameter would start to decrease. As a rough initial estimate, following the saltatory conduction hypothesis, one could take , where is the membrane capacitance of the myelinated section and is the capacitance of the node of Ranvier – in such a case if we take, for example 1 µ node of Ranvier, 250 µ long myelinated section, axon radius of 5 µ and thickness of the myelin 2.5 µ the corresponding would be around 0.5. Using similar parameters as in the classical paper by Hodgkin and Huxley [23] (see Table 1), a numerical example using parameter value (assuming 1 µ nodes of Ranvier, 50 µ length for myelinated sections and taking for the numerical example) and [µm] follows (see Fig. 11). Choosing (which is an upper limit case) means that we are making here an assumption that saltatory conduction velocity is much faster than signal propagation velocity in an unmyelinated axon. We have to emphasize here that the “unit cell” we are modelling in the computational node contains both the node of Ranvier and the myelinated section adjacent to it, i.e., the parameters are a mix of “pure” Lieberstein model for unmyelinated axon and myelination effects like in [3].
@@
The AP propagation velocity as a function of the -ratio as a function of axon radius (in the [µm] range) is depicted in Fig. 12. The case corresponds to an unmyelinated axon. It should be noted that here we are assuming that the thickness of the myelin (related to g-ratio, which is included through coefficient (18)) is the same even if -ratio is varied. The upper limit case () for the conduction velocity is used. It should be noted that the dependence of the AP velocity on the axon radius is still there even in the myelinated case. The larger the radius of the axon, the faster the velocity of the AP increases as the -ratio is increasing.
@@ 11 Discussion and summary
We started with the model for unmyelinated axon (13), (14) [31, 49] in a form which is similar to the Maxwell equations for a transmission line and used the work of Lieberstein as a source of inspiration for handling the question of inductance in the context of signal propagation along the nerve axon. The key difference between the classical HH model and the Lieberstein-inspired model used here is that the mechanism for signal propagation along the axon emerges like a wave as a consequence of opting to keep the inductivity . While, indeed, there exist variations of the classical HH model which support AP signal propagation where the equations are written in the form of PDE instead of the usual ODE form (which describes signal evolution in time at a fixed spatial point). In these, normally, the potential-gradient-type member responsible for propagating the signal along the axon is not as clearly defined from the viewpoint of physics (usually some kind of abstracted diffusion-type process is used).
@@
We briefly explained the saltatory conduction mechanism hypothesis as we prepared to modify the governing equations to take into account the effect of myelination on the propagation of the AP along the axon. The model (13), (14) based on Maxwell equations for a transmission line is then modified to include the influence of myelination. In addition to the g-ratio normally considered in the earlier studies (taken into account indirectly through coefficient ) which takes into account the myelination geometry perpendicular to the axon we introduce the so-called “myelination-ratio” or -ratio which describes the influence of myelin distribution on the signal propagation in the direction of the axis of the axon. The numerical example (see Fig. 12), using parameters from the earlier studies, demonstrates physiologically plausible behaviour for the model. The model is reduced to the description of the unmyelinated axon if the length of the myelinated sections along the axon is taken as zero. Under the considered parameter combinations we can observe the AP propagation velocities up to 79 (the signal propagation velocity range for myelinated axons is given as roughly 10 to 100 in the earlier studies [33, 44]). For comparison, the model yields for the unmyelinated () axon in the same parameter range propagation velocity of about 7.8 at [µm].
@@
It is important to emphasize that the proposed continuum-based model is philosophically similar to how the transmission line equations are composed. The ‘unit-cell’ in the context of the myelinated axon in the model is composed of the node of Ranvier and the myelinated section next to it. This is opposed to the alternative approach where the processes in nodes of Ranvier are described by the classical HH model while myelinated sections are handled separately either through some numerical scheme or by an alternative model coupled with the HH equations in the node of Ranvier through some mechanism. Having a relatively simple pair of PDEs which are connected to the Maxwell equations for anything involving the movement of charges in an environment could be considered superior to investigating causal connections and making predictions than something that is not as clearly connected to the basic physical considerations.
@@ Acknowledgements
This research was supported by the Estonian Research Council (PRG 1227). Jüri Engelbrecht acknowledges the support from the Estonian Academy of Sciences. The authors thank Prof. Antoine Jerusalem from the University of Oxford for helpful comments during the preparation of the present paper.
Author contributions
Kert Tamm contributed by writing manuscript text and composing figures, performing numerical simulations and analysis of the results. Tanel Peets contributed analysis for comparing the modified Lieberstein equation against the classical Hodgkin-Huxley equations. All other work (like novel ideas presented, physical interpretation of the results, proofreading, etc) is a collaborative effort equally contributed by all authors.